# Anomalous charge density wave in two-dimensional altermagnet WO

### v1

@@ Acknowledgements.
This work was financially supported by the National Natural Science Foundation of China (Grant No.12434009, No.12204533 and No.62476278), the National Key RD Program of China (Grant No. 2024YFA1408601), the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China (Grant No. 24XNKJ15).Computational resources have been provided by the Physical Laboratory of High Performance Computing at Renmin University of China.References
@@
- González-Hernández et al. [2021] R. González-Hernández, L. mejkal, K. Výborný, Y. Yahagi, J. Sinova, T. c. v. Jungwirth, and J. elezný, Efficient Electrical Spin Splitter Based on Nonrelativistic Collinear Antiferromagnetism, Phys. Rev. Lett. 126, 127701 (2021).
@@
- Hwang et al. [2024] J. Hwang, W. Ruan, Y. Chen, S. Tang, M. F. Crommie, Z.-X. Shen, and S.-K. Mo, Charge density waves in two-dimensional transition metal dichalcogenides, Rep. Prog. Phys. 87, 044502 (2024).
- Sødequist and Olsen [2024] J. Sødequist and T. Olsen, Two-dimensional altermagnets from high throughput computational screening: Symmetry requirements, chiral magnons, and spin-orbit effects, Appl. Phys. Lett. 124, 182409 (2024).
- Bai et al. [2024] L. Bai, W. Feng, S. Liu, L. Šmejkal, Y. Mokrousov, and Y. Yao, Altermagnetism: Exploring new frontiers in magnetism and spintronics, Adv. Funct. Mater. 34, 2409327 (2024).
@@
- Wu, Bin and Song, Yong-liang and Ji, Wei-xiao and Wang, Pei-ji and Zhang, Shu-feng and Zhang, Chang-wen [2023] Wu, Bin and Song, Yong-liang and Ji, Wei-xiao and Wang, Pei-ji and Zhang, Shu-feng and Zhang, Chang-wen, Quantum anomalous Hall effect in an antiferromagnetic monolayer of MoO, Phys. Rev. B 107, 214419 (2023).

### v2

@@ Acknowledgements.
We thank Jun-Wei Liu for valuable discussions. This work was financially supported by the National Natural Science Foundation of China (Grant No.12434009, No.12204533, No.12304165 and No.62476278), the National Key RD Program of China (Grant No. 2024YFA1408601), the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China (Grant No. 24XNKJ15). Computational resources have been provided by the Physical Laboratory of High Performance Computing at Renmin University of China. Z.-H. D. and L. W. contributed equally to this work.References
@@
- González-Hernández et al. [2021] R. González-Hernández, L. mejkal, K. Výborný, Y. Yahagi, J. Sinova, T. Jungwirth, and J. elezný, Efficient Electrical Spin Splitter Based on Nonrelativistic Collinear Antiferromagnetism, Phys. Rev. Lett. 126, 127701 (2021).
@@
- Chen et al. [2016] C.-W. Chen, J. Choe, and E. Morosan, Charge density waves in strongly correlated electron systems, Rep. Prog. Phys. 79, 084505 (2016).
- Agterberg et al. [2020] D. F. Agterberg, J. S. Davis, S. D. Edkins, E. Fradkin, D. J. Van Harlingen, S. A. Kivelson, P. A. Lee, L. Radzihovsky, J. M. Tranquada, and Y. Wang, The physics of pair-density waves: cuprate superconductors and beyond, Annu. Rev. Condens. Matter Phys. 11, 231 (2020).
@@
- Regmi et al. [2025] R. B. Regmi, H. Bhandari, B. Thapa, Y. Hao, N. Sharma, J. McKenzie, X. Chen, A. Nayak, M. El Gazzah, B. G. Márkus, et al., Altermagnetism in the layered intercalated transition metal dichalcogenide CoNb4Se8, Nat. Commun. 16, 4399 (2025).
@@
- Wu et al. [2023] B. Wu, Y.-L. Song, W.-X. Ji, P.-J. Wang, S.-F. Zhang, and C.-W. Zhang, Quantum anomalous Hall effect in an antiferromagnetic monolayer of MoO, Phys. Rev. B 107, 214419 (2023).
@@
- [53] See supplemental material at http://link.aps.org/xxx, which includes a detailed description of computational methods as well as supplemental figures.
- Giannozzi et al. [2009] P. Giannozzi, S. Baroni, N. Bonini, M. Calandra, R. Car, C. Cavazzoni, D. Ceresoli, G. L. Chiarotti, M. Cococcioni, I. Dabo, et al., Quantum espresso: a modular and open-source software project for quantumsimulations of materials, J. Phys. Condens. Matter 21, 395502 (2009).
@@
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- Togo [2023] A. Togo, First-principles phonon calculations with phonopy and phono3py, J. Phys. Soc. Jpn. 92, 012001 (2023).
- Momma and Izumi [2011] K. Momma and F. Izumi, VESTA3 for three-dimensional visualization of crystal, volumetric and morphology data, J. Appl. Crystallogr. 44, 1272 (2011).
- Ganose et al. [2021] A. M. Ganose, A. Searle, A. Jain, and S. M. Griffin, Ifermi: A python library for fermi surface generation and analysis, J. Open Source Softw. 6, 3089 (2021).
- Hamann [2013] D. R. Hamann, Optimized norm-conserving vanderbilt pseudopotentials, Phys. Rev. B 88, 085117 (2013).
- Baroni et al. [2001] S. Baroni, S. de Gironcoli, A. Dal Corso, and P. Giannozzi, Phonons and related crystal properties from density-functional perturbation theory, Rev. Mod. Phys. 73, 515 (2001).